# Tensor Product Attention Is All You Need

### v1

@@ 3.3 KV Caching and Memory Reduction
For large $h$ and $d_{h}$ (typically $d_{h}=64$ or $128$), setting $R_{K},R_{V}\ll d_{h}$ (e.g., $1$ or $2$) often yields 10×10$\times$10 × or more reduction.
Table 1: Comparison of different attention mechanisms. Here, $R_{Q}$, $R_{K}$, and $R_{V}$ denote the ranks for queries, keys, and values in TPA, respectively. Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in Section 3.5. For MLA, $d_{h}^{R}$ and $d_{h}$ are the dimensions for RoPE and non-RoPE parts; $d_{c}^{\prime}$ and $d_{c}$ are the dimensions of compressed vectors for query and key-value, respectively.
@@ 3.4.1 MHA as Non-contextual TPA
[…]
    so that $\mathbf{e}_{i}\otimes\cdot$ corresponds to the $i$-th head of $\mathbf{Q}_{t}$.
@@
MQA and GQA as Non-Contextual TPA. Multi-Query Attention (MQA) (Shazeer, 2019) and Grouped Query Attention (GQA) (Ainslie et al., 2023) also emerge naturally from TPA by restricting the head-dimension factors to be non-contextual and low-rank:
@@
    $\mathbf{K}_{t}=(1,\ldots,1)^{\top}\;\otimes\;\mathbf{b}^{K}(\mathbf{x}_{t}),\quad\mathbf{V}_{t}=(1,\ldots,1)^{\top}\;\otimes\;\mathbf{b}^{V}(\mathbf{x}_{t}),$
@@
3.5 Other Variants of TPA
and similarly for keys/values. This arrangement is effective if the token-dimension structure remains mostly uniform across the sequence, while the head-dimension factors capture context.
TPA KV Only. One can preserve a standard query mapping,
$\mathbf{Q}_{t}=\bm{W}^{Q}\,\mathbf{x}_{t}\in\mathbb{R}^{h\times d_{h}},$
$\mathbf{b}^{K}_{r}(\mathbf{x}_{t})=\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),$
@@ 3.6 Model Architectures
We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention). Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times h\times d_{h}}$ defined in Section 3.1, T6 utilize the overall architecture of LLaMA (Touvron et al., 2023) while changing the self-attention block to our TPA-based version. The feed-forward network (FFN) adopts a SwiGLU layer, as in (Shazeer, 2020; Touvron et al., 2023).
@@
Rotary Positional Embedding (RoPE). As discussed in Section 3.2, RoPE (Su et al., 2024b) is applied to the $\mathbf{Q}$ and $\mathbf{K}$. Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see (3.6) and Theorem 1.
@@
SwiGLU Feed-Forward Network. Following Shazeer (2020); Touvron et al. (2023), our T6 uses a SwiGLU-based Feed-Forward Network (FFN):
$\operatorname{FFN}(\mathbf{x})=\bigl{[}\sigma(\mathbf{x}\,\bm{W}_{1})\,\odot\,(\mathbf{x}\,\bm{W}_{2})\bigr{]}\,\bm{W}_{3},$
where $\sigma$ is the SiLU (a.k.a., swish) nonlinearity, $\odot$ is element-wise product, and $\bm{W}_{1},\bm{W}_{2},\bm{W}_{3}$ are learnable parameters. Note that other activation functions can also be used.
@@ 4.1 Language Modeling Tasks
All experiments reported in this paper are implemented on the nanoGPT code base (Karpathy, 2022), using the FineWeb-Edu 100B dataset (Lozhkov et al., 2024). The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation. We compare T6 against the baseline Llama architecture (Touvron et al., 2023) with SwiGLU activation (Shazeer, 2020) and RoPE embeddings (Su et al., 2024a), as well as Llama variants that replace Multi-Head Attention (MHA; Vaswani et al., 2017) with Multi-Query Attention (MQA; Shazeer, 2019), Grouped Query Attention (GQA; Ainslie et al., 2023), or Multi-head Latent Attention (MLA; Liu et al., 2024a). In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer. We train models at four scales: small (124M parameters), medium (353M), and large (773M). Details on architecture hyperparameters and training hardware appear in Appendix B.1.
@@
Training Setup. We follow the nanoGPT training configuration. In particular, we use the AdamW (Loshchilov, 2017) optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$. We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler (Loshchilov & Hutter, 2016) with $2{,}000$ warmup steps and a (total) global batch size of $480$. For the small, medium, and large models, we set maximum learning rates of $6\times 10^{-4}$, $3\times 10^{-4}$, and $2\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $3\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
@@
Training & Validation Curves. Figures 2 and 3 compare training and validation loss curves for the large (773M) and medium (353M) models on FineWeb-Edu-100B. Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final losses. For instance, in Figure 2(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages. Meanwhile, Multi-Head Latent Attention (MLA) (Liu et al., 2024a) (blue curves) generally trains more slowly and yields higher losses.
@@
Validation Perplexity. Figure 4 shows the validation perplexities of the medium- and large-scale models. Mirroring the loss curves, TPA and TPA-KVonly steadily outperform MHA, MQA, GQA, and MLA over the course of training. By the end of pretraining (around $49$B tokens), TPA-based approaches achieve the lowest perplexities in most configurations.
@@
Downstream Evaluation. We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (Yadav et al., 2019), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2020) and MMLU (Hendrycks et al., 2021), using the lm-evaluation-harness codebase (Gao et al., 2024). For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy. Tables 89 in the appendix present results for small models; Tables 23 for medium models; Tables 45 for large models;
@@
For the medium-size (353M) models (Tables 23), TPA generally ties or outperforms all competing methods, achieving, for example, an average of 51.41% in zero-shot mode versus MHA’s 50.11%, MQA’s 50.44%, and MLA’s 48.96%. When given two-shot prompts, TPA again leads with 53.12% average accuracy. A similar trend appears for the large-size (773M) models (Tables 45), where TPA-KVonly attains the highest average (53.52% zero-shot, 55.33% two-shot), closely followed by full TPA.
@@
(a) Training Loss
(b) Validation Loss
Figure 3: The training loss and validation loss of medium-size (353M) models with different attention mechanisms on the FineWeb-Edu 100B dataset.
(a) Validation Perplexity of Medium Models
(b) Validation Perplexity of Large Models
Figure 4: The validation perplexity of medium-size (353M) models and large-size (773M) models with different attention mechanisms on the FineWeb-Edu 100B dataset.
Table 2: The evaluation results of medium models with different attention mechanisms pretrained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
@@
MHA 56.52 29.27 58.84 44.06 35.00 68.44 51.07 25.35 76.40 49.44
MQA 55.68 28.24 60.86 44.17 35.20 68.66 52.72 25.14 72.90 49.29
GQA 54.88 29.61 56.36 43.77 35.20 68.82 52.57 25.41 74.80 49.05
MLA 55.30 29.27 58.96 41.92 35.40 67.25 51.78 25.20 75.60 48.96
TPA 59.30 31.91 60.98 45.57 34.60 69.48 53.91 24.93 77.20 50.88
Table 3: The evaluation results of medium models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
MHA 64.44 32.85 59.05 44.18 33.20 68.72 50.12 26.01 87.40 49.44
MQA 64.27 32.94 57.71 44.36 31.80 68.01 51.70 25.99 86.00 49.29
GQA 61.70 32.17 52.81 43.99 33.80 68.50 53.35 24.44 86.40 50.80
MLA 62.75 30.80 59.17 42.02 34.80 67.08 52.41 26.11 84.80 51.10
TPA 66.54 34.47 58.96 45.35 33.00 69.21 53.99 24.51 91.30 53.04
@@
Method ARC-E ARC-C BoolQ HellaSwag OBQA PIQA WG MMLU SciQ Avg.
MHA 67.85 36.35 59.82 50.22 35.00 70.67 53.35 23.92 91.10 54.25
MQA 68.86 36.09 53.79 50.50 37.00 70.89 54.70 25.01 88.00 53.87
GQA 69.15 36.09 58.84 50.29 36.20 70.73 54.22 26.08 90.00 54.62
MLA 68.56 35.41 60.12 49.18 38.00 69.21 55.25 25.29 88.20 54.36
TPA 70.41 37.71 60.06 51.30 34.00 71.06 54.54 25.79 90.30 55.02
@@ 5 Related Work
Transformers and Attention. As a sequence-to-sequence architecture Transformer (Vaswani et al., 2017) introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies. Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns (Child et al., 2019; Shi et al., 2023; Han et al., 2024; Liang et al., 2024a; Li et al., 2024; Liang et al., 2024b), kernel-based projections (Choromanski et al., 2021), and linearized transformers (Tsai et al., 2019; Katharopoulos et al., 2020; Schlag et al., 2021; Zhang et al., 2023b; Sun et al., 2023; Zhang et al., 2024). To decrease memory usage and circumvent the limitation of memory bandwidth in training, Shazeer (2019) proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head. To tackle with the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) (Ainslie et al., 2023) divides queries into several groups, and each group of queries shares a single key head and value head. Recently, DeepSeek-V2 (Liu et al., 2024a) applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value. In comparison to the approaches above, TPA applied a low-rank tensor product to compute the queries, keys, and values where the cached representations of keys and values are much smaller than those in MHA, achieving better reduction on memory assumption of KV cache in inference time.
KV Cache Optimization. During the inference time of Transformers, key and value tensors of the previous tokens are repeatedly computed due to their auto-regressive nature. To enhance efficiency, firstly proposed by Ott et al. (2019), these tensors can be cached in memory for future decoding, referred to as the KV cache. However, the KV cache requires additional memory usage and may add to more latencies due to the bandwidth limitation (Adnan et al., 2024). Therefore, previous studies have explored diverse approaches to mitigate these issues, including KV cache eviction to discard less significant tokens (Zhang et al., 2023c; Xiao et al., 2024; Cai et al., 2024; Adnan et al., 2024), dynamic sparse attention among selected keys and values (Ribar et al., 2024; Tang et al., 2024; Singhania et al., 2024), KV cache offloading to CPU (He & Zhai, 2024; Lee et al., 2024; Sun et al., 2024), as well as quantization of KV cache (Xiao et al., 2023; Liu et al., 2024c; Hooper et al., 2024). Besides these methods, it is also effective to reduce the amount of KV cache for each token, by approaches such as reducing the number of KV heads (Ren et al., 2024; Ainslie et al., 2023), cross-layer KV re-usage (Xiao et al., 2019; Mu et al., 2024; Wu et al., 2024), and low-rank KV representation (Saxena et al., 2024). Different from the methods above, TPA reduces the size of the KV cache by using tensor-decomposed keys and values.

### v2

@@ 3.3 KV Caching and Memory Reduction
For large $h$ and $d_{h}$ (typically $d_{h}=64$ or $128$), setting $R_{K},R_{V}\ll h$ (e.g., rank $1$ or $2$) often yields 10×10$\times$10 × or more reduction.
Table 1: Comparison of different attention mechanisms. Here, $R_{Q}$, $R_{K}$, and $R_{V}$ denote the ranks for queries, keys, and values in TPA, respectively. Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in Section F. For MLA, $d_{h}^{R}$ and $d_{h}$ are the dimensions for RoPE and non-RoPE parts; $d_{c}^{\prime}$ and $d_{c}$ are the dimensions of compressed vectors for query and key-value, respectively.
@@ 3.4.1 MHA as Non-contextual TPA
    $\mathbf{a}^{Q}_{i}=R_{Q}\mathbf{e}_{i}\in\mathbb{R}^{h},$ (3.8)
    where $\mathbf{e}_{i}\in\mathbb{R}^{h}$ is the $i$-th standard basis vector, so that $\mathbf{e}_{i}\otimes\cdot$ corresponds to the $i$-th head of $\mathbf{Q}_{t}$.
@@
MQA and GQA as Non-Contextual TPA. Multi-Query Attention (MQA) (Shazeer, 2019) and Grouped Query Attention (GQA) (Ainslie et al., 2023) [^6] The original definitions of MQA and GQA are presented in Appendix E.1 and E.2, respectively. also emerge naturally from TPA by restricting the head-dimension factors to be non-contextual and low-rank:
@@
    $\mathbf{K}_{t}$ $=(1,\ldots,1)^{\top}\;\otimes\;\mathbf{b}^{K}(\mathbf{x}_{t}),$
    $\mathbf{V}_{t}$ $=(1,\ldots,1)^{\top}\;\otimes\;\mathbf{b}^{V}(\mathbf{x}_{t}),$
@@
3.5 Computational Cost.
For a detailed analysis of the computational cost of TPA, please refer to Appendix A, which shows that the training and inference flops of TPA with optimized implementation (without materializing $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$) are smaller than MHA, GQA, and MLA. Specifically, when we set $R_{q}=6$, $R_{k}=R_{v}=2$ (our default setting), TPA is 10×10$\times$10 × or more faster on calculating $\mathbf{Q}\mathbf{K}^{\top}$ than MLA during inference (see Appendix A.8).
@@ 3.6 Model Architectures
We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention). Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilize the overall architecture of LLaMA (Touvron et al., 2023) while changing the self-attention block to our TPA-based version. The feed-forward network (FFN) adopts a SwiGLU layer, as in (Shazeer, 2020; Touvron et al., 2023).
@@
Rotary Positional Embedding (RoPE). As discussed in Section 3.2, RoPE (Su et al., 2024) is applied to the $\mathbf{Q}$ and $\mathbf{K}$. Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see (3.6) and Theorem 1.
@@
SwiGLU Feed-Forward Network. Following Shazeer (2020); Touvron et al. (2023), our T6 uses a SwiGLU-based Feed-Forward Network (FFN): $\operatorname{FFN}(\mathbf{x})=\bigl{[}\sigma(\mathbf{x}\,\bm{W}_{1})\,\odot\,(\mathbf{x}\,\bm{W}_{2})\bigr{]}\,\bm{W}_{3}$, where $\sigma$ is the SiLU (a.k.a., swish) nonlinearity, $\odot$ is element-wise product, and $\bm{W}_{1},\bm{W}_{2},\bm{W}_{3}$ are learnable parameters. Note that other activation functions can also be used.
@@ 4.1 Language Modeling Tasks
All experiments reported in this paper are implemented on the nanoGPT code base (Karpathy, 2022), using the FineWeb-Edu 100B dataset (Lozhkov et al., 2024). The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation. We compare T6 against the baseline Llama architecture (Touvron et al., 2023) with SwiGLU activation (Shazeer, 2020) and RoPE embeddings (Su et al., 2024), as well as Llama variants that replace Multi-Head Attention (MHA; Vaswani et al., 2017) with Multi-Query Attention (MQA; Shazeer, 2019), Grouped Query Attention (GQA; Ainslie et al., 2023), or Multi-head Latent Attention (MLA; Liu et al., 2024a). In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer. We train models at four scales: small (124M parameters), medium (353M), large (773M), and XL (1.5B). Details on architecture hyperparameters and training hardware are shown in Appendix G.1.
@@
Training Setup. We follow the nanoGPT training configuration. In particular, we use the AdamW (Loshchilov, 2017) optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$. We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler (Loshchilov & Hutter, 2016) with $2{,}000$ warmup steps and a (total) global batch size of $480$. For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $3\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $6\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
@@
Training & Validation Curves. Figures 2 and 3 compare training and validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B. Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see F) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final losses. For instance, in Figure 3(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages. Meanwhile, Multi-Head Latent Attention (MLA) (Liu et al., 2024a) (blue curves) generally trains more slowly and yields higher losses.
@@
Validation Perplexity. Figure 4 (in the Appendix) shows the validation perplexities of the medium- and large-scale models. Mirroring the loss curves, TPA and TPA-KVonly steadily outperform MHA, MQA, GQA, and MLA over the course of training. By the end of pretraining (around $49$B tokens), TPA-based approaches achieve the lowest perplexities in most configurations.
@@
Downstream Evaluation. We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (Yadav et al., 2019), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2020) and MMLU (Hendrycks et al., 2021), using the lm-evaluation-harness codebase (Gao et al., 2024). For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy. Due to the page limitation, we only display the zero-shot evaluation results of medium and large models here in Tables 2 and 3. Zero-shot evaluation of small and XL models are displayed in Tables 6 and 7 in the appendix. Moreover, we also present 2-shot evaluation results in Tables 8910 and 11 in the appendix.
@@
For the medium-size (353M) models (Tables 2 and 9), TPA generally ties or outperforms all competing methods, achieving, for example, an average of 51.41% in zero-shot mode versus MHA’s 50.11%, MQA’s 50.44%, and MLA’s 50.13%. When given two-shot prompts, TPA again leads with 53.12% average accuracy. A similar trend appears for the large-size (773M) models (Tables 3), where TPA-KVonly attains the highest average (53.52% zero-shot). And for the XL size (1.5B) models (Table 7), TPA-KVonly attains the highest average (55.03% zero-shot).
@@
(a) Medium models (353M)
(b) Large models (773M)
(c) XL models (1.5B)
Figure 2: The training loss of medium-size (353M), large-size (773M) as well as XL-size (1.5B) models, with different attention mechanisms on the FineWeb-Edu 100B dataset.
(a) Medium models (353M)
(b) Large models (773M)
(c) XL models (1.5B)
Figure 3: The validation loss of medium-size (353M), large-size (773M) as well as XL-size (1.5B) models, with different attention mechanisms on the FineWeb-Edu 100B dataset.
Table 2: The evaluation results of medium models with different attention mechanisms pre-trained using FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
@@
MHA 59.51 29.52 59.60 45.68 34.20 68.82 53.43 23.33 76.90 50.11
MQA 57.62 31.91 59.45 45.69 35.40 69.31 53.51 26.47 74.60 50.44
GQA 58.67 31.48 58.29 45.45 35.20 68.50 54.46 24.58 76.50 50.35
MLA 56.65 29.52 57.83 46.05 34.60 69.42 52.80 24.62 79.70 50.13
TPA (non-ctx-A) 58.96 31.48 59.76 45.07 34.80 69.21 53.59 25.42 76.40 50.52
TPA 58.38 31.57 59.39 46.83 37.00 70.02 54.06 25.52 79.90 51.41
@@ 5 Related Work
Transformers and Attention. As a sequence-to-sequence architecture Transformer (Vaswani et al., 2017) introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies. Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns (Child et al., 2019; Shi et al., 2023; Han et al., 2024; Liang et al., 2024a; Li et al., 2024; Liang et al., 2024b), kernel-based projections (Choromanski et al., 2021), and linearized transformers (Tsai et al., 2019; Katharopoulos et al., 2020; Schlag et al., 2021; Zhang et al., 2023b; Sun et al., 2023; Zhang et al., 2024). To decrease memory usage and circumvent the limitation of memory bandwidth in training, Shazeer (2019) proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head. To tackle with the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) (Ainslie et al., 2023) divides queries into several groups, and each group of queries shares a single key head and value head. Recently, DeepSeek-V2 (Liu et al., 2024a) applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value. Concurrently, Hu et al. (2024) proposed Multi-matrix Factorization Attention (MFA), which can be simply seen as MQA with low-rank factorized Q. Compared to the approaches above, TPA applied contextual tensor decompositions to represent queries, keys, and values activations compactly, achieving better reduction on the size of KV cache with improved performance.
KV Cache Optimization. During the inference time of Transformers, key and value tensors of the previous tokens are repeatedly computed due to their auto-regressive nature. To enhance efficiency, firstly proposed by Ott et al. (2019), these tensors can be cached in memory for future decoding, referred to as the KV cache. However, the KV cache requires additional memory usage and may add to more latencies due to the bandwidth limitation (Adnan et al., 2024). Therefore, previous studies have explored diverse approaches to mitigate these issues, including KV cache eviction to discard less significant tokens (Zhang et al., 2023c; Xiao et al., 2024; Cai et al., 2024; Adnan et al., 2024), dynamic sparse attention among selected keys and values (Ribar et al., 2024; Tang et al., 2024; Singhania et al., 2024), KV cache offloading to CPU (He & Zhai, 2024; Lee et al., 2024; Sun et al., 2024), as well as quantization of KV cache (Xiao et al., 2023; Liu et al., 2024c; Hooper et al., 2024). Different from the methods above, TPA reduces the size of the KV cache by using tensor-decomposed KV.